Solve by completing the square:
w^2 - 2w- 15 = 0

Solve
x^2 + 2x - 8 < or = to 0

Answer by checkley71(8403)

W~2-2W=15 OR W~2-2W+1=15+1 OR (W-1)(W-1)=16 OR W-1=4 OR W=5
X~2+2X-8=0 OR (X+4)(X-2)=0 OR X=-4 & X=+2